The following data are obtained from the records of a factory


Sales (40,000 units @ Br. 25 each) ………… Br. 100,000


Less: Variable Costs………………………… 70,000


Contribution Margin ……………………….. 30,000


Less: Fixed Costs …………………………… 18,000


Net Profit …………………………………… 12.000


Required: Compute:


1) The number of units by selling which the company will neither loss or gain anything


2) The sales needed to earn a profit of 20% on sales


3) The extra units which should be sold to obtain the present profit if it is proposed to


reduce the selling price by 25%


4) The selling price to be fixed to bring down its break-even point to Br.500 units under


present condition

1
Expert's answer
2022-01-18T09:59:12-0500

1) Break Even point= "\frac {Fixed cost}{(Sales per unit price-Variable cost per unit)}"

="\frac{18000}{(25-1.75)}=\frac{18000}{23.25}"

=774.2 units

2) Let us assume sales = X

Profit will be 20% of X= 0.2X

Profit =sales -Cost

Cost = Fixed cost + Variable cost

=18000+70,000= 88,000

0.2X=X-88,000

X-0.2X=88,000

0.8X=88,000

X= 110,000 units


3) Extra Units to be sold

Current Selling Price=75% 0f Br25= Be 18.75

Profit = Sales- cost

Profit remains at 12,000

Let sales be X

12000=X-88,000

X=100000

Number of units sold="\frac {total sales}{selling price per unit}"

="\frac{100000}{18.75}=5333.3"

Extra units= 5333.3-4000= 1,333.33

4) Selling price to be fixed to bring break even to Br 500 units

Break even="\frac {Fixed cost}{(Sales per unit price-Variable cost per unit)}"

Let Selling price per unit be X

500="\frac {18000}{(X-1.75)}"

"500\times(X-1.75)=18000"

"500X-875=18000"

500X=18875

X="\frac{18875}{500}=37.75"